STEINER CORPORATION

Partial Balance Sheet

December 31, 2018

Stockholders’ equity

       8% Preferred stock, $100 par value, cumulative, 5,000 shares issued ˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑ$500,000

       In excess of par value— preferred stock ˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑ    $280,000

       Common stock, $5 par value, 440,000 shares issued ˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑ$2,200,000
       In excess of par value—common stock ˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑ      $800,000           

       Total paid-in capital ˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑ    $3,780,000

           

       Retained earnings ˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑ $1,334,000                                                       

       Total paid-in capital and retained earnings ˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑ$5,114,000

            Less: Treasury stock (10,000 common shares) ˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑ($120,000)

       Total stockholders’ equity ˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑˑ $4,994,000

On December 1, 2018, the board of directors declared cash dividends. This is the only cash dividend declared during 2018. The dividend in arrears as of December 1, 2018 is $30,000. The net income was $600,000 for the entire year of 2018. The balance of retained earnings at the end of 2017 was $900,000. Compute the amount of cash dividend on common stock declared on December 1, 2018 in .

On January 12, 2019, Steiner issued 20,000 shares of common stock with an issue price of $20 per share, $5 par value. Record the journal entry for this transaction in .

On January 28, 2019, Steiner reissued 5,000 shares of its treasury stock for $15 per share. Record the journal entry for this transaction in .

On March 28, 2019, Steiner reissued 4,000 shares of its treasury stock for $10 per share. Record the journal entry for this transaction in .

On December 20, 2019, Steiner declared a 10% stock dividend to common shareholders. The market value of common stock was $25 per share at the declaration of the stock dividend. Enter the value of the stock dividend to common shareholders in .
Answer #1
Steiner Corporation                                                            -  
1-Dec-18 Cash Dividends Declared          166,000 =900,000+600,000-1,334,000
Preferred Dividends in Arrears          (30,000) (Op RE + Net Income - Clsg RE)
Common Dividends Declared          136,000
Date Account Debit Credit Calculation
12-Jan-19 Cash          400,000 =20,000*20
Common Stock          100,000 =20,000*5
Paid in Capital in Excess of Par-Common Stock          300,000
28-Jan-19 Cash            75,000 =5,000*15
Treasury Stock            60,000 =5,000*(120,000/10,000)
Paid in Capital from Sale of Treasury Stock            15,000
28-Mar-19 Cash            40,000 =4,000*10
Paid in Capital from Sale of Treasury Stock              8,000
Treasury Stock            48,000 =4,000*(120,000/10,000)
20-Dec-19 Retained Earnings      1,147,500 =45,900*25
Common Stock Dividends Distributable          229,500 =45,900*5
Paid in Capital in Excess of Par-Common Stock          918,000
Workings
Stock dividends
Opening Shares          440,000
Shares Issued            20,000
Number of Common Shares Issued          460,000
Opening Shares - Treasury Stock          (10,000)
Treasury Stock reissued              5,000
Treasury Stock reissued              4,000
Number of Common Shares O/s          459,000
10% stock dividend 10%
Stock dividends            45,900 =459,000*10%